OpenStudy (anonymous):

Determine the input that yields the minimum value for the function f(x)=square root of (x^2-8x+41) I just don't know how to find the input when it is the square root of

OpenStudy (raden):

y = sqrt(x^2-8x+41) = sqrt((x-4)^2 + 25) to get the minimum value, it should when x=4

OpenStudy (raden):

so, just put x=4 to that function what is the value of y will u get ?

OpenStudy (raden):

y=f(x)

OpenStudy (raden):

any way u can use derivative to get it, but i think thought one like on

OpenStudy (calculusfunctions):

Do you know calculus?